OpenStudy (anonymous):

Kyle ran 2/3 mi on Monday, 1 1/6 mi on Tuesday, and 7/12 mi on Wednesday. How many total miles did Kyle run? Express the answer in simplest form

OpenStudy (anonymous):

you have to add

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

make all the bottoms of the fraction the same

OpenStudy (anonymous):

I got...

OpenStudy (anonymous):

\[1\frac{ 5 }{ 12 }\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ADD 2/3+ 1 1/6, which is 1 5/12+7/12= 1/1=1 ANSWER

OpenStudy (anonymous):

can you help me @Tyler.Evan2

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with what @1903

OpenStudy (anonymous):

my question

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

nvm the answer is \[2\frac{ 5 }{ 12 }\]